OpenStudy (samigupta8):

if the straight line through the point P(3,4) make an angle of π/6 with the x-axis and meets the line 12x+5y+10=0 at Q find the length of PQ

OpenStudy (shubhamsrg):

what is the eqn of this line passing through P ?

OpenStudy (samigupta8):

that we cn find out

OpenStudy (samigupta8):

by using point slope form \[y-4=\frac{ 1 }{\sqrt{3} }(x-3)\]

OpenStudy (shubhamsrg):

right, find the intersection point of this line and 12x +5y +10=0 find that point's distance from (3,4) and voila! you have an answer

OpenStudy (samigupta8):

hey i m getting value of y as\[y=\frac{ 48\sqrt{3}-46 }{ 12\sqrt{3}+5 }\]

OpenStudy (samigupta8):

this is so complicated.........

OpenStudy (shubhamsrg):

your y should be -2(36 sqrt3 -89)/37

OpenStudy (samigupta8):

how u getting dis value im not getting any such value

OpenStudy (samigupta8):

what is ur final answer in this question